# RingMoE: Mixture-of-Modality-Experts Multi-Modal Foundation Models for Universal Remote Sensing Image Interpretation

### v1

@@ 3.4.2 Dense Expert Integration Schemes
Knowledge Compressing. Beyond the two aforementioned schemes of direct expert knowledge aggregation, we aim to preserve key knowledge from each expert while ensuring effective compression. Low-rank decomposition is a powerful technique for model compression, enabling the transformation of a high-dimensional matrix into a low-rank representation while maximally retaining its information. To achieve this, we compress each expert’s weight matrix into a low-rank form [47] and subsequently merge them into a high-rank matrix with the same dimensionality as the original, as illustrated in Fig.7.
@@
To enhance usability across various downstream tasks, the pruned and modularized models offer flexible deployment options. Modality-specific sub-models can be employed individually for unimodal tasks or integrated for multi-modal scenarios. Additionally, users can choose among full fine-tuning for peak accuracy, parameter freezing for computational efficiency, or parameter-efficient fine-tuning (PEFT) to achieve an optimal balance between accuracy and resource usage (see Appendix Fig.A3.b). Such designs enable the proposed model to be deployed flexibly across geographically distributed computing resources, which conforms to the emerging requirements of collaboration between different scales of computing platforms. In this case, appropriately selecting the platform, modal-specialized sub-models, and model scales according to the RS application requirements can significantly improve the resource utilization and the sustainability of the RSFMs, highlighting RingMoE’s versatility for real-world applications.
@@ 3.5 RingMOSS: A Comprehensive Multi-Modal Pre-Training Dataset
To establish a versatile RSFM adaptable to diverse tasks, RingMoE requires a comprehensive and diverse pre-training dataset that accounts for significant variations in imaging modalities, spatial resolutions, temporal dynamics, geographic regions, and scene complexities. To meet this challenge, we curate RingMOSS, a large-scale multi-modal RS dataset comprising 400 million images from nine satellite platforms, covering a broad spectrum of Earth observation scenarios.
RingMOSS spans four key modalities: MS, Opt, and SAR in complex-valued and amplitude forms (i.e., SAR-L1, and SAR-L2) (see Appendix Fig.A1-A2). Opt imagery constitutes the largest portion of the dataset, with 302,862,240 images, representing 70.1% of the total dataset. MS data includes 35,931,400 images, accounting for 20.2%. SAR data, sourced from the GF-3 satellite, is available in two forms: SAR-L1, in complex-valued with amplitude and phase information, and SAR-L2, containing only amplitude information. By integrating amplitude and phase information, SAR provides reliable, all-weather coverage, capturing surface contours, elevation, and structural details, thus overcoming the inherent limitations of Opt and MS data [42, 43]. Further details on the dataset, including collection strategies and pre-processing, can be found in Appendix B.1.
@@ 4 Experiments
This section provides a comprehensive evaluation of RingMoE, emphasizing its practical performance. Sec.4.1 evaluates RingMoE on the challenging task of recognizing unseen classes using only a single labeled sample, without additional fine-tuning, across five benchmarks. Then, Sec.4.2 compares RingMoE with SOTA foundation and task-specific models across 20 public benchmarks, covering six distinct RS interpretation tasks. Sec.4.3 presents ablation studies and in-depth analyses. Notably, the full 14.7B multi-modal model is decomposed into multiple 6.5B single-modal sub-models for evaluation, enabling both unimodal and multi-modal applications. Additionally, we assess pruned variants, including RingMoE-EP (4.3B for Opt), RingMoE-KC (1B), RingMoE-KA (1B), and RingMoE-KS (1B), providing insights into scalability and resource-efficient deployment.
Further implementation details of RingMoE, including pre-training settings (Appendix B.3) and downstream tasks such as dataset descriptions, experimental settings, and evaluation metrics (Appendix B.4), are provided in the Appendix. Due to space constraints, the main body primarily focuses on comparative analysis against the top two to three methods for each dataset, visualized through bar charts and radar plots. For a more comprehensive perspective, Appendix C presents detailed quantitative and qualitative comparisons.
@@ 4.1 Performance on Unseen Class Recognition
Due to the rapid changes in RS task requirements, comprehensive labeling of RS data has become impractical. To address this challenge, models need the ability to generalize to unseen classes with a few labeled examples, enabling faster and more flexible interpretation of novel data. To this end, we directly apply the RingMoE model to few-shot tasks (including few-shot classification [58] and few-shot segmentation [59]) to evaluate its generalization capability. These tasks aim to quickly generalize to unseen classes with only a few labeled samples, without focusing on previously trained classes. Notably, RingMoE serves as the image encoder for feature extraction in these tasks, with its parameters fixed during both the training and testing phases (See Appendix B.4.8 for detailed settings).
Few-shot classification experiments (Fig.9 and Appendix Tab.3) on the NWPU-RESIS45 [60], WHU-R19 [61], and UCM [62] datasets reveal that RingMoE with 6.5B parameters achieves the highest accuracies: 78.91%, 85.25%, and 66.50%. Even after compression to 1B parameters using expert pruning techniques such as knowledge summing and averaging, e.g., RingMoE-KS and RingMoE-KA, our models retain SOTA performance. In few-shot segmentation (Fig.9 and Appendix Tab.4), similar trends are observed on iSAID [63] and LoveDA [64], where the 6.5B parameter model achieves 52.45% and 32.40% mIoU, respectively. Even the pruned 4.3B model (i.e., RingMoE-EP) scores 51.78% and 30.96% mIoU, outperforming current methods. Importantly, the 1B parameter version still maintains competitive SOTA performance, underscoring the model’s powerful feature extraction capabilities.
The above results highlight RingMoE’s ability to generalize effectively with minimal labeled data while maintaining strong performance even with reduced model sizes (down to 1B parameters). Its balance of accuracy and data efficiency makes it well-suited for real-world tasks, enabling rapid adaptation to novel, unseen data in resource-constrained environments with limited annotated samples.
Figure 10: Feature visualization comparison across RSFMs on Scene Classification tasks. We extract features from various RSFMs on the test sets of both classification datasets and employ t-SNE to visualize the feature distributions of 10 randomly selected categories, highlighting the feature clustering across different models.
@@ 4.2 Performance on Single/Multi-modal Interpretation Tasks
We validate RingMoE across 20 benchmarks spanning single and multi-modal tasks, covering diverse data modalities and common tasks in remote sensing. In scene classification, we evaluate the performance of various parameter-scale versions of RingMoE. For other downstream tasks like segmentation and detection, we focus on the 1B version of RingMoE, considering practical applications and deployment constraints. Notably, this section centers on the analysis of performance outcomes, while details on datasets and experimental settings for each task are provided in Appendix B.4.
(1) Scene classification. We evaluate RingMoE and other RSFMs on two popular optical classification datasets, NWPU-RESISC45 [60] and AID [65] (see Tab.II). RingMoE demonstrates strong performance with limited data, achieving 95.90% and 98.19% accuracies on NWPU-RESISC45 and AID, respectively, surpassing SOTA SkySense [27] by 1.05% and 0.51%. In addition, we extract the features from various foundation models on the test set of the two datasets and randomly select 10 categories for visual representation using t-SNE (see Fig.10). Results indicate that RingMoE exhibits a more discriminative representation capability than other foundation models, characterized by tighter intra-class clustering and larger inter-class variance.
To support low-cost deployment, we explore several pruning strategies to compress the model. The 4.3B-parameter RingMoE-EP (Opt), employing sparse expert pruning, achieves 95.43% and 97.88% accuracies, closely matching the full model. Remarkably, the 1B-parameter version, compressed through dense expert integration, retains high classification accuracy, surpassing other foundation models and performing comparably to the 2B-parameter SkySense. These findings underscore RingMoE’s robust feature extraction capabilities, even with significant parameter reductions. Given its practical advantages, we further evaluate the 1B-parameter version (RingMoE-KC) across other downstream tasks.
@@
  • Optical/Multi-spectral segmentation. We evaluate RingMoE and other RSFMs on two optical datasets, iSAID [63] and Potsdam, as well as the multi-spectral dataset Dyna.-pla [71], using mIoU and mF1 as metrics (Tab.III and Fig.11.a). RingMoE achieves SOTA performance on the Dyna.-pla dataset and ranks second on iSAID and Potsdam, just behind SkySense [27]. Notably, RingMoE accomplishes these results with a compressed 1B-parameter version, highlighting its efficiency and effectiveness. Confusion matrix analysis in Appendix Fig.A4 confirms RingMoE’s stable segmentation performance and ability to distinguish categories effectively. Qualitative results in Fig.11.b and Appendix Fig.A6.a-b highlight its superior detail capture, particularly for features like rivers and roads, outperforming comparable RSFMs like Satlas [68] and SatMAE [20].
  • SAR segmentation. We also evaluate RingMoE and other specialized methods on two SAR datasets: the complex-valued SAR dataset SARSegL1 [72] and the amplitude SAR dataset AIR-POLSAR-SEG [73] (Fig.11.a). On SARSegL1, RingMoE achieves significant improvements, surpassing existing methods in mIoU and OA by 18.21% and 9.91%, respectively, with the highest IoU across all categories (Appendix Tab.5). On AIR-POLSAR-SEG, it surpasses the second-best method by 5.33% in mIoU, with notable gains in Industrial Area (12.8%), Natural Area (4.19%), Water (6.35%), and Housing (3.35%) (Appendix Tab.6). Qualitative results in Fig.11.b and Appendix Fig.A6.c-d highlight its superior precision in processing complex SAR features.
  • Multi-modal segmentation. RingMoE is evaluated on two multi-modal segmentation datasets, WHU-OPT-SAR [74] and DFC23 [75], which combine optical and amplitude SAR modalities, using mIoU and OA metrics (Fig.11.a and Appendix Tab.7-8). It achieves SOTA performance on both datasets, with mIoU and OA of 54.7% and 84.1% on WHU-OPT-SAR, and 73.4% and 95.8% on DFC23. Fig.11.b and Appendix Fig.A6.e showcases segmentation examples where RingMoE demonstrates precise delineation of village and road areas on WHU-OPT-SAR and reduces roof category confusion with finer boundaries on DFC23, highlighting its advanced multi-modal segmentation capabilities.
@@
  • Optical detection. RingMoE is evaluated on three optical detection datasets: DIOR-horizontal (DIOR) [77], DIOR-rotation (DIOR-R), and HRSC2016 [78] (Tab.IV and Fig.12.a). It achieves SOTA performance on DIOR and DIOR-R, surpassing SkySense [27] by 3.72% and 1.77% mAP50, respectively. On HRSC2016, RingMoE sets a new benchmark, exceeding the previous SOTA by 1.5% mAP50. These results are further supported by qualitative visualizations in Fig.12.b and Appendix Fig.A7.a, underscoring its superior optical detection capabilities.
  • SAR Detection. RingMoE achieves state-of-the-art performance on two amplitude SAR detection datasets, with 94.2% mAP50 on HRSID [79] (+1.2%) for ship detection and an 8.1% mAP50 improvement on SAR-AIRcraft-1.0 [80] for plane detection (Fig.12.a and Appendix Tab.9). Fig.12.b and Appendix Fig.A7.b highlights RingMoE’s robustness in complex scenarios such as harbors and airports, showcasing its adaptability to SAR detection tasks.
@@
(4) Object tracking. RingMoE achieves SOTA performance on the AIR-MOT [81] and AIR-HSAO [82] optical tracking datasets (Fig.12.a and Appendix Fig.A5.b). On AIR-MOT, it improves MOTA by 1.3% and IDF1 by 1.9% over the previous SOTA. Similarly, on AIR-HSAO, RingMoE achieves a 2.2% improvement in MOTA and a 2.0% increase in IDF1. Sample visualizations in Fig.12.c and Appendix Fig.A8.a demonstrate RingMoE’s capability to accurately capture and track objects in complex remote sensing scenarios.
@@
Figure 12: Performance Comparison across Object Detection, Object Tracking, Change Detection, and Depth Estimation tasks. a. Quantitative Analysis: RingMoE is quantitatively compared with current state-of-the-art (SOTA) methods across four tasks. Object detection is evaluated using mAP50, object tracking is assessed with MOTA and IDF1, change detection uses F1 score and IoU, while depth estimation is measured with Rel, $\delta{1}$, $\delta_{2}$, and $\delta_{3}$ as primary indicators. b,c,d,e. Qualitative Analysis: Selected examples from detection, tracking, change detection, and depth estimation tasks highlight RingMoE’s adaptability and superior performance across diverse scenarios.
@@
(5) Change detection. Fig.12.a and Appendix Tab.10 showcase RingMoE’s superior performance on the LEVIR-CD [83] and CDD [84] datasets, outperforming all competing RSFMs and task-specific methods to achieve the highest metrics on both benchmarks. By leveraging its pre-trained foundation architecture, RingMoE effectively extracts generalized features, significantly improving semantic change detection in bi-temporal images. Visualizations (Fig.12.d and Appendix Fig.A8.b) illustrate its accuracy in detecting changes with minimal false positives, setting new standards for remote sensing change detection.
@@
(6) Depth estimation. RingMoE achieves SOTA performance on two large-scale remote sensing depth estimation benchmarks (Fig.12.a and Appendix Tab.11). On the Vaihingen dataset, RingMoE reduces the Rel error by 35% compared to the previous leading model, Heightformer, and significantly improves accuracy metrics $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ by 15.9%, 0.5%, and 1.0%, respectively, marking a substantial advancement in depth estimation. Similarly, on the Potsdam dataset, RingMoE achieves the best performance. Visualization results on the ISPRS Vaihingen dataset (Fig.12.e and Appendix Fig.A8.c) further validate RingMoE’s precision, closely aligning with ground truth, particularly in areas dense with buildings and trees. These findings highlight RingMoE’s capability for accurate and detailed depth estimation.

### v2

@@ 3.4.2 Dense Expert Integration Schemes
Knowledge summing (KS) and averaging (KA). For Knowledge Summing, given two linear projections $W_{k;1}^{*}\in\mathbb{R}{{}^{d_{1}\times d_{2}}}$ and $W_{k;2}^{*}\in\mathbb{R}{{}^{d_{2}\times d_{1}}}$ in every expert (modal-specialized, cooperative, and shared) in the RMoE layer, we sum the weights of all experts to aggregate their knowledge. For Knowledge Averaging, similar to summing, we average the weights of the experts.
Knowledge Compressing. Beyond direct aggregation, we propose a low-rank fusion strategy to retain essential knowledge from all experts while achieving effective compression. As illustrated in Fig. 7, each expert’s weight matrix is approximated utilizing Singular Value Decomposition (SVD):
@@
Another projection $W_{2}\in\mathbb{R}{{}^{d_{2}\times d_{1}}}$ has a similar process.
To enhance adaptability across downstream tasks, the pruned and modularized RingMoE supports flexible deployment. Modality-specific sub-models can operate independently for unimodal tasks or be combined for multimodal applications. Users may select from full fine-tuning, parameter freezing, or parameter-efficient fine-tuning (PEFT) to balance accuracy and efficiency (see Appendix Fig.3.b). Such flexibility enables deployment across heterogeneous and geographically distributed computing platforms, improving resource utilization and sustainability in practical RS applications.
@@ 3.5 RingMOSS: A Comprehensive Multi-Modal Pre-Training Dataset
To establish a versatile RSFM adaptable to diverse tasks, RingMoE requires a comprehensive and diverse pre-training dataset that accounts for significant variations in imaging modalities, spatial resolutions, temporal dynamics, geographic regions, and scene complexities. Thus, we curate RingMOSS [^2] All public data, along with detailed documentation are available at: https://github.com/HanboBizl/RingMoEDatasets., a large-scale multi-modal RS dataset comprising 400 million images from nine satellite platforms, covering a broad spectrum of Earth observation scenarios.
RingMOSS spans four key modalities: MS, Opt, and SAR in complex-valued and amplitude forms (i.e., SAR-L1, and SAR-L2) (see Appendix Fig.1-2). Opt imagery constitutes the largest portion of the dataset, with 302,862,240 images, representing 70.1% of the total dataset. MS data includes 35,931,400 images, accounting for 20.2%. SAR data, sourced from the GF-3 satellite, is available in two forms: SAR-L1, in complex-valued with amplitude and phase information, and SAR-L2, containing only amplitude information. By integrating amplitude and phase information, SAR provides reliable, all-weather coverage, capturing surface contours, elevation, and structural details, thus overcoming the inherent limitations of Opt and MS data [40, 20]. Further details on the dataset, including collection strategies and pre-processing, can be found in Appendix B.1.
@@ 4 Experiments
This section provides a comprehensive evaluation of RingMoE, emphasizing its practical performance. Sec.4.1 evaluates RingMoE on the challenging task of recognizing unseen classes using only a single labeled sample, without additional fine-tuning, across five benchmarks. Then, Sec.4.2 compares RingMoE with SOTA foundation and task-specific models across 20 public benchmarks, covering six distinct RS interpretation tasks. Sec.4.3 presents ablation studies and in-depth analyses. Notably, the full 14.7B multi-modal model is decomposed into multiple 6.5B single-modal sub-models for evaluation, enabling both unimodal and multi-modal applications. Additionally, we assess pruned variants, including RingMoE-EP (4.3B for Opt), RingMoE-KC (1B), RingMoE-KA (1B), and RingMoE-KS (1B), providing insights into scalability and resource-efficient deployment (See Appendix E.1 for more comprehensive evaluation).
Further implementation details of RingMoE, including pre-training settings (Appendix B.3) and downstream tasks such as dataset descriptions, experimental settings, and evaluation metrics (Appendix B.5), are provided in the Appendix. For comprehensive performance comparisons, Appendix C presents both quantitative and qualitative analyses, while Appendix E offers additional ablation studies on expert pruning strategies, SAR-L1 pretraining effects, and the impact of the physics-informed loss used in SAR-L1.
@@ 4.1 Performance on Unseen Class Recognition
Due to the rapid changes in RS task requirements, comprehensive labeling of RS data has become impractical. To address this challenge, models need the ability to generalize to unseen classes with a few labeled examples, enabling faster and more flexible interpretation of novel data. To this end, we directly apply the RingMoE model to few-shot tasks (including few-shot classification [56] and few-shot segmentation [57]) to evaluate its generalization capability. These tasks aim to quickly generalize to unseen classes with only a few labeled samples. Notably, RingMoE serves as the image encoder for feature extraction in these tasks, with its parameters fixed during both the training and testing phases (See Appendix B.5.8 for detailed settings).
Few-shot classification experiments (Fig.9 and Appendix Tab.4) on the NWPU-RESIS45 [58], WHU-R19 [59], and UCM [60] datasets reveal that RingMoE with 6.5B parameters achieves the highest accuracies: 78.91%, 85.25%, and 66.50%. Even after compression to 1B parameters using expert pruning techniques such as knowledge summing and averaging, e.g., RingMoE-KS and RingMoE-KA, our models retain SOTA performance. In few-shot segmentation (Fig.9 and Appendix Tab.5), similar trends are observed on iSAID [61] and LoveDA [62], where the 6.5B parameter model achieves 52.45% and 32.40% mIoU, respectively. Even the pruned 4.3B model (i.e., RingMoE-EP) scores 51.78% and 30.96% mIoU, outperforming current methods. Importantly, the 1B parameter version still maintains competitive SOTA performance, significantly outperforming general foundation models such as SAM [63], VRP-SAM [64], underscoring the model’s powerful feature extraction capabilities.
The above results highlight RingMoE’s robust generalization with minimal labeled data while maintaining leading performance even with reduced model sizes (down to 1B parameters). It offers a practical solution for real-world RS tasks under limited labeling and resource constraints.
@@ 4.2 Performance on Single/Multi-modal Interpretation Tasks
We validate RingMoE across 20 benchmarks spanning single and multi-modal tasks, covering diverse data modalities and common tasks in remote sensing. We also compare RingMoE with recent multimodal large language models (MLLMs) to highlight its advancements. In scene classification, we evaluate the performance of various parameter-scale versions of RingMoE. For other downstream tasks like segmentation and detection, we focus on the 1B version of RingMoE, considering practical applications and deployment constraints. Detailed experimental settings for each task are provided in Appendix B.5.
(1) Scene classification. We evaluate RingMoE and other RSFMs on two optical classification benchmarks, NWPU-RESISC45 [58] and AID [65] (Tab.II). RingMoE achieves 95.90% and 98.19% accuracies, exceeding the previous SOTA SkySense [26] by 1.05% and 0.51%, respectively. T-SNE visualizations (Appendix Fig.4) further confirm its discriminative feature representation, showing tighter intra-class clusters and larger inter-class separation than competing RSFMs. For efficient deployment, we investigate pruning-based compression. The 4.3B RingMoE-EP (Opt), using sparse expert pruning, achieves 95.43% and 97.88% accuracies, nearly matching the full model. Even the 1B dense-compressed variant (RingMoE-KC) maintains strong performance, surpassing most RSFMs and approaching the 2B SkySense. These results demonstrate that RingMoE preserves powerful feature extraction and generalization ability even under significant parameter reduction. Given its practical advantages, we further evaluate the 1B-parameter version (RingMoE-KC) across other downstream tasks.
@@
  • Optical/Multi-spectral segmentation. We evaluate RingMoE and other RSFMs on two optical datasets, iSAID [61] and Potsdam [71], as well as the multi-spectral dataset Dyna.-pla [72] (Tab.III and Fig.10.a). RingMoE achieves SOTA performance on the Dyna.-pla dataset and ranks second on iSAID and Potsdam, just behind SkySense [26]. Notably, RingMoE accomplishes these results with a compressed 1B-parameter version, highlighting its efficiency and effectiveness. Confusion matrix analysis in Appendix Fig.5 reveals RingMoE’s stable per-category performance. Qualitative results in Fig.10.b and Appendix Fig.7.a-b highlight its superior detail capture, outperforming comparable RSFMs like Satlas [68] and SatMAE [19].
  • SAR segmentation. We evaluate RingMoE and several specialized methods on two SAR datasets: the complex-valued SAR-L1 dataset SARSegL1 [73] and the amplitude SAR-L2 dataset AIR-POLSAR-SEG [74] (Fig.10.a). On SARSegL1, RingMoE achieves significant improvements, surpassing existing methods in mIoU and OA by 18.21% and 9.91%, respectively, with the highest IoU across all categories (Appendix Tab.6). On AIR-POLSAR-SEG, it surpasses the second-best method by 5.33% in mIoU, with notable gains in Industrial Area (12.8%), Natural Area (4.19%) and Water (6.35%) (Appendix Tab.7). Qualitative results in Fig.10.b and Appendix Fig.7.c-d highlight its superior precision in processing complex SAR features.
  • Multi-modal segmentation. RingMoE is evaluated on two multi-modal segmentation datasets, WHU-OPT-SAR [75] and DFC23 [76], which combine optical and amplitude SAR-L2 modalities, using mIoU and OA metrics (Fig.10.a and Appendix Tab.8-9). It achieves SOTA performance on both datasets, with mIoU and OA of 54.7% and 84.1% on WHU-OPT-SAR, and 73.4% and 95.8% on DFC23. Fig.10.b and Appendix Fig.7.e qualitatively demonstrate RingMoE’s advanced multimodal segmentation capabilities.
@@
  • Optical detection. RingMoE is evaluated on three optical detection datasets: DIOR-horizontal (DIOR) [78], DIOR-rotation (DIOR-R), and HRSC2016 [79] (Tab.IV and Fig.11.a). It achieves SOTA performance on DIOR and DIOR-R, surpassing SkySense [26] by 3.72% and 1.77% mAP50, respectively. On HRSC2016, RingMoE also exceeded the previous SOTA by 1.5% mAP50. These results are further supported by qualitative visualizations in Fig.11.b and Appendix Fig.8.a, underscoring its superior optical detection capabilities.
  • SAR Detection. RingMoE achieves SOTA performance on two amplitude SAR-L2 detection datasets, with 94.2% mAP50 on HRSID [80] (+1.2%) for ship detection and an 8.1% mAP50 improvement on SAR-AIRcraft-1.0 [81] for plane detection (Fig.11.a and Appendix Tab.10). Fig.11.b and Appendix Fig.8.b highlights its robustness in complex scenarios such as harbors and airports, showcasing its adaptability to SAR detection.
@@
(4) Object tracking. RingMoE achieves SOTA performance on the AIR-MOT [82] and AIR-HSAO [83] optical tracking datasets (Fig.11.a and Appendix Fig.6.b). On AIR-MOT, it improves MOTA by 1.3% and IDF1 by 1.9% over the previous SOTA. Similarly, on AIR-HSAO, RingMoE achieves a 2.2% improvement in MOTA and a 2.0% increase in IDF1. Sample visualizations in Fig.11.c and Appendix Fig.9.a demonstrate RingMoE’s capability to accurately capture and track objects in complex remote sensing scenarios.
@@
Figure 11: Performance Comparison across Object Detection, Object Tracking, Change Detection, and Depth Estimation tasks. a. Quantitative Analysis: RingMoE is quantitatively compared with current SOTA methods across four tasks. On the left column, object detection (red labels) is evaluated using mAP50, object tracking (blue labels) is assessed with MOTA and IDF1. On the right column, change detection (red labels) uses F1 score and IoU, while depth estimation (blue labels) is measured with Rel, $\delta{1}$, $\delta_{2}$, and $\delta_{3}$ as primary indicators. b,c,d,e. Qualitative Analysis: Selected examples from detection, tracking, change detection, and depth estimation tasks highlight RingMoE’s adaptability and superior performance across diverse scenarios.
@@
(5) Change detection. Fig.11.a and Appendix Tab.11 showcase RingMoE’s superior performance on the LEVIR-CD [84] and CDD [85] datasets, outperforming all competing RSFMs and task-specific methods to achieve the highest metrics on both benchmarks. By leveraging its pre-trained foundation architecture, RingMoE effectively extracts generalized features, significantly improving semantic change detection in bi-temporal images. Visualizations (Fig.11.d and Appendix Fig.9.b) illustrate its accuracy in detecting changes with minimal false positives.
@@
(6) Depth estimation. RingMoE achieves SOTA performance on two large-scale RS depth estimation benchmarks, ISPRS Vaihingen and Potsdam [71] (Fig.11.a and Appendix Tab.12). On the Vaihingen dataset, RingMoE reduces the Rel error by 35% compared to the previous leading model, Heightformer, and significantly improves accuracy metrics $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$ by 15.9%, 0.5%, and 1.0%, respectively, marking a substantial advancement in depth estimation. Similarly, on the Potsdam dataset, RingMoE achieves the best performance. Visualization results on the ISPRS Vaihingen dataset (Fig.11.e and Appendix Fig.9.c) further validate RingMoE’s precision, closely aligning with ground truth, particularly in areas dense with buildings and trees. These findings highlight RingMoE’s capability for accurate and detailed depth estimation.
(7) Comparison with MLLMs. Recent multimodal large language models (MLLMs) have shown strong generalization in vision-language tasks. We compare RingMoE with general MLLMs (e.g., MiniGPTv2 [6], LLaVA-1.5 [86]) and RS-specific MLLMs (e.g., GeoChat [87], EarthGPT [88]) on several key RS tasks: optical classification, optical/SAR-L2 detection (Appendix Tab.13–15 in D). RingMoE consistently outperforms all MLLMs. On NWPU-RESISC45 classification, it achieves 96.28% OA using only 20% of training data. On DIOR detection, it surpasses the best RS-MLLMs by over 17% mAP50. Notably, on the challenging SARDet-100k [89], RingMoE reaches 89.72% mAP50, while general MiniGPTv2 completely fail (0.00%). These results demonstrate the effectiveness of RingMoE in perception and semantic interpretation across diverse remote sensing scenarios.